There were 3 Nos. of 10 Ton each A/C Plant Make M/s. Blue Star Ltd., which were in use since 1978. These Plants were then replaced by 3 x 10 ton KIRLOSKAR Make AC Plant. Due to very frequent failure of compressors, these plants did not work satisfactorily till July 2000. The plants were re-commissioned with bigger compressors (QR 15) and modified filters on 14.08.2000.. The scale guard was also installed in the water line of the system in Yr. 2001.The compressor QRK 15 of plant No. 3 failed on 23.05.2006 and is replaced by 10 Ton semi hermetic bock compressor (dated 29.02.08) under PRG funds of CE(WZ). The compressor of Unit No.2 failed in Nov.2008. Since the Unit had already served a useful time of 10 years, it was replaced by 11 ton Air cooled condensing unit under IEBR(METP) 2008-09 of make Blue Star, Model No.DSA 1321 on 03.08.2008. Later other remaining 2 plants were also replaced by 11 ton Blue Star make Air cooled units. Presently all these 3 units are working satisfactorily.
AC Plant Transmitting Centre
At Transmitting centre, 3 Nos. of 18 Ton each compressor along with 2 AHU were working since October 1987 by M/s. Batliboi & co., Mumbai. Under the replacement scheme of old 100 KW MW X’TR BEL- Make , 4 x 15 Ton Make - M/s. Snowcool system & Engg., Nagpur was provided. . All the old 3 x 10 Ton Batliboi & Co. Bombay – Make Plants were dismantled on- Feb. 2006 for the installation of these new 4 x 15 Ton air cooled plants. A new plenum chamber as per the requirement of A/C Plant was fabricated and installed. New supply ducts were laid in front side of the X’TR Hall, emergency Play back studio and testing room.
The false ceiling itself has been used as a return duct for these rooms and brought to return chamber. A New power supply cable has been laid from L.T. Panel to A/C Plant room. Out door units of A/C Plants was installed on the back side of plant room on the platform and covered with AC sheet shed to protect from rain & sun rays. The installations of A/C Plants were completed by Dec. 2007.
The Receiving Centre at Rawanbhata has been declared obsolete in 1993. All the equipments and aerials have been dismantled & kept in Transmitting centre. NonExchange line between transmitter and R/C as well as Telephone was surrendered to P & T Department. All the three electricity service connections were also surrendered to M.P.E.B. in June 1999. Disposal of 11.52 Acres of R/C land and building is under consideration with Directorate.
The agreement between A.I.R Raipur & VSNL, Mumbai for the purpose of execution of communication line for audio program & data exchange was made on 17.04.2007. 128 Kbps ISDN line for INMARSAT terminal at A.I.R., Raipur was procured on April 2007. The S/I/T/C was carried out by M/s. WEBEL Mediatronics on 28.01.2008 and later on the service Engineer of the said firm imparted formal training to Tech staff of A.I.R Raipur.
Water Supply Transmitter
A new Bore-well has been digged in the X’TR premises and submersible pump has been installed to uplift the water upto existing 10,000 Ltrs capacity Over-head tank in addition to old bore well with compressor type pump assembly.
At Transmitting centre, two Nos. of 11 KV feeders are terminated at 6-pole structure in A.I.R. campus along with metering facility, from Chhattisgarh State Electricity board. Under 10th Plant and replacement of 100 KW MW X’TR following works were carried out :
1. Two length of H.T. cable from 6 pole structure to H.T. switch Gear room replaced.
2. Old 02 Nos H.T. O.C.Bs replaced by new 4 panel H.T. VCB.
3. 02 Nos. 11 KV / 450 volts H.T. Transformer have been replaced by new one.
4. Earlier, there were no A/B switches at 6 pole structure. For maintenance purpose, 02 Nos. A/B switches are installed.
5. L.T. switchgear panel has been replaced by new one with provision of essential and non-essential light and power load. The Output of 125 KVA D/G set is also terminated on this panel.
6. All the old Kit Kat type I/C switches in distribution panel were replaced by MCB for Light and power of old building.
7. The electrical light and power wiring along with switch and socket has been replaced by new one in old building.
8. Twin light decorative tube rod fittings have been replaced by mirror optic twin light tube rod fittings in X’TR Hall and Emergency studio.
9. All the other single and double tube rod fittings in old building, ATU and pump room have been replaced by new one.
10. Flood lights above the main building and ATU has been replaced by new one.
11. Street light fitting along with poles have been replaced by new one.
The old D/G set 62.5 KVA, Make – Ruston Sl. No. 48G1178 has been taken out of the circuit in Feb. 2007 and diverted to AIR Ambikapur & new D/G set 125 KVA provided by M/s. Jeevan Diesels electrical Division, Silvasa has been installed in Feb. 2007. During the year HT cable of MPSRTC feeder became defective, hence it was replaced with the new one, now both the feeders are working satisfactorily. The HT feeder line namely MPSRTC feeder cable was damaged and it was repaired twice but it could not work for long time. Insulation was found poor, hence the same was replaced with a new 11 KV HT cable.
Power Supply Studios
There were two independent 400 V feeders for Studio-cum-office complex earlier. The voltage on one of the feeder (locally called LIC feeder) remained always low. Due to this, A/C Plants were unable to be put on this particular feeder. This problem was discussed with officials of C.S.EB. They suggested to obtain the H.T. connection in lieu of existing LT ones as in the mean time, 1 KW FM X’TR, 10 KW IGNOU X’TR, CES etc. also came into existence . Hence, H.T. connection of 11 KV with contract Demand 178 KVA was energized on 19.08.2006. In year 1998, CE (WZ) under MR Scheme supplied a new 25 KVA D/G set. It was installed and brought in operation on 22.04.1998. The old 22.5 KVA D/G set was auctioned in 2003. Another new DG set of 62.5 KVA , Kirloskar make is provided for 10 KW FM transmitter. The old power supply cables of studio and office complex were replaced by new ones in March 2009, and LT panels 02 nos. of switch gear room were replaced by new panel in March 2010.